OpenStudy (anonymous):

Can someone check my work? I will fan and medal 1. Solve the equation. 2x – 3 = 15 x = 6 x = 9 <<< x = 24 x = 36 2. Solve the equation. r/3 (r over 3) – 4 = 11 r = 45 r = 35 <<< I’m not really sure about this one r = 33 r = 28 Solve the equation. 7 + 3m = 28 m = 18 <<<< m = 11 2/3 m = 9 m = 7

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first one is right' you can check via \[2\time s9-3=15\]

OpenStudy (misty1212):

\[\frac{35}{3}-4=11\] is unlkely

OpenStudy (anonymous):

Yea. I doubted #2 was right

OpenStudy (misty1212):

\[\frac{x}{3}-4=11\] add 4\[\frac{x}{3}=15\] multiply by 3 \[x=45\]

OpenStudy (misty1212):

last one is wrong too ;(

OpenStudy (misty1212):

a) subtract 7 b) divide by 3

OpenStudy (anonymous):

Ah well. Number two makes sense now, so thanks. :)

OpenStudy (misty1212):

ok good let me know what you get for #3

OpenStudy (anonymous):

I got m=7 for number 3

OpenStudy (anonymous):

Yeah. I think that's the answer

OpenStudy (anonymous):

Yeah. I think that's the answer

OpenStudy (misty1212):

yeah it is 1

OpenStudy (misty1212):

\[7+3\times 7=7+21=28\] so 7 for sure

OpenStudy (anonymous):

Yay! Well, thanks for helping! :3

OpenStudy (misty1212):

\[\huge \color\magenta\heartsuit\]